### v1

# RapVerse: Coherent Vocals and Whole-Body Motions Generations from Text

Abstract
In this work, we introduce a challenging task for simultaneously generating 3D holistic body motions and singing vocals directly from textual lyrics inputs, advancing beyond existing works that typically address these two modalities in isolation. To facilitate this, we first collect the RapVerse dataset, a large dataset containing synchronous rapping vocals, lyrics, and high-quality 3D holistic body meshes. With the RapVerse dataset, we investigate the extent to which scaling autoregressive multimodal transformers across language, audio, and motion can enhance the coherent and realistic generation of vocals and whole-body human motions. For modality unification, a vector-quantized variational autoencoder is employed to encode whole-body motion sequences into discrete motion tokens, while a vocal-to-unit model is leveraged to obtain quantized audio tokens preserving content, prosodic information and singer identity. By jointly performing transformer modeling on these three modalities in a unified way, our framework ensures a seamless and realistic blend of vocals and human motions. Extensive experiments demonstrate that our unified generation framework not only produces coherent and realistic singing vocals alongside human motions directly from textual inputs, but also rivals the performance of specialized single-modality generation systems, establishing new benchmarks for joint vocal-motion generation. The project page is available for research purposes at https://vis-www.cs.umass.edu/RapVerse.
1 Introduction
Figure 1: RapVerse. We present a unified text-conditioned multi-modality generation framework, for jointly generating holistic body motions and singing vocals from textual lyrics inputs only. Note that the corresponding video frames are just shown for reference.
In the evolving landscape of multi-modal content generation in terms of sound and motion, significant strides have been made in individual modalities, including text-to-music schneider2023mo ; agostinelli2023musiclm ; huang2023noise2music , text-to-vocal liu2022diffsinger , text-to-motion guo2022generating ; zhang2023t2m ; chen2023executing ; jiang2024motiongpt ; lu2023humantomato , and audio-to-motion yi2023generating ; habibie2021learning ; liu2023emage generation. These developments have paved the way for creating more dynamic and interactive digital content. Despite these advancements, existing works predominantly operate in silos, addressing each modality in isolation. However, there’s strong psychological evidence that for human beings, the generation of sound and motion are highly related and coupled lakoff1980metaphors . A unified system for joint generation allows for a more expressive and nuanced communication of emotions, intentions, and context, where the generation of one modality could guide and assist the other in a coherent and efficient way.
In this paper, we tackle a crucial problem: can a machine not only sing with emotional depth but also perform with human-like expressions and motions? We propose a novel task for generating coherent singing vocals and whole-body human motions (including body motions, hand gestures, and facial expressions) simultaneously, see Fig. 1. This endeavor holds practical significance in fostering more immersive and naturalistic digital interactions, thereby elevating virtual performances, interactive gaming, and the realism of virtual avatars.
An important question naturally arises: what constitutes a good model for unified generation of sound and motion? Firstly, we consider textual lyrics as the proper form of inputs for the unified system, since text provides a highly expressive, interpretable flexible means of conveying information by human beings, and could serve as a bridge between various modalities. Previous efforts explore scores liu2022diffsinger , action commands zhang2023t2m ; chen2023executing ; jiang2024motiongpt , or audio signals yi2023generating as inputs, which are inferior to textual inputs in terms of semantic richness, expressiveness and flexible integration of different modalities.
Secondly, we reckon that a joint generation system that could produce multi-modal outputs simultaneously is better than a cascaded system that executes the single-modal generation sequentially. A cascaded system, combining a text-to-vocal module with a vocal-to-motion module, risks accumulating errors across each stage of generation. For instance, a misinterpretation in the text-to-vocal phase can lead to inaccurate motion generation, thereby diluting the intended coherence of the output. Furthermore, cascaded architectures necessitate multiple training and inference phases across different models, substantially increasing computational demands.
To build such a joint generation system, the primary challenges include: 1) the scarcity of datasets that provide lyrics, vocals, and 3D whole-body motion annotations simultaneously; and 2) the need for a unified architecture capable of coherently synthesizing vocals and motions from text. In response to these challenges, we have curated RapVerse, a large-scale dataset featuring a comprehensive collection of lyrics, singing vocals, and 3D whole-body motions. Despite the existence of datasets available for text-to-vocal liu2022diffsinger ; huang2021multi ; duan2013nus ; sharma2021nhss , text-to-motion plappert2016kit ; mahmood2019amass ; guo2022generating ; lin2024motion , and audio-to-motion cao2017realtime ; habibie2021learning ; ginosar2019learning ; fanelli20103 ; cudeiro2019capture ; wuu2022multiface , the landscape lacks a unified dataset that encapsulates singing vocals, whole-body motion, and lyrics simultaneously. Most notably, large text-to-vocal datasets huang2021multi ; zhang2022m4singer are predominantly in Chinese, limiting their applicability for English language research and lacking any motion data. And text-to-motion datasets plappert2016kit ; guo2022generating ; lin2024motion typically focus on text descriptions of specific actions paired with corresponding motions without audio data, often not covering whole body movements. Moreover, audio-to-motion datasets liu2022diffsinger ; lu2023co focus primarily on speech rather than singing. A comparison of existing related datasets is demonstrated in Table. 1. The RapVerse dataset is divided into two distinctive parts to cater to a broad range of research needs: 1) a Rap-Vocal subset containing a large number of pairs of vocals and lyrics, and 2) a Rap-Motion subset encompassing vocals, lyrics, and human motions. The Rap-Vocal subset contains 108.44 hours of high-quality English singing voice in the rap genre without background music. Paired lyrics and vocals are crawled from the Internet from 32 singers, with careful cleaning and post-processing. On the other hand, the Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters pavlakos2019expressive using the annotation pipeline of Motion-X lin2024motion , synchronous singing vocals and corresponding lyrics.
With the RapVerse dataset, we explore how far we can push by simply scaling autoregressive multimodal transformers with language, audio, and motion for a coherent and realistic generation of vocals and whole-body human motions. To this end, we unify different modalities as token representations. Specifically, three VQVAEs van2017neural are utilized to compress whole-body motion sequences into three-level discrete tokens for head, body, and hand, respectively. For vocal generation, previous works min2021meta ; donahue2020end ; liu2022diffsinger ; min2021meta share a common paradigm, producing mel-spectrograms of audio signals from input textual features and additional music score information, following with a vocoder oord2016wavenet ; valin2019lpcnet ; yang2021multi to reconstruct the phase. We draw inspiration from the speech resynthesis domain polyak2021speech , and learn a self-supervised discrete representation to quantize raw audio signal into discrete tokens while preserving the vocal content and prosodic information. Then, with all the inputs in discrete representations, we leverage a transformer to predict the discrete codes of audio and motion in an autoregressive fashion. Extensive experiments demonstrate that this straightforward unified generation framework not only produces realistic singing vocals alongside human motions directly from textual inputs but also rivals the performance of specialized single-modality generation systems.
To sum up, this paper makes the following contributions:
  • We release RapVerse, a large dataset featuring synchronous singing vocals, lyrics, and high-quality 3D holistic SMPL-X parameters.
  • We design a simple but effective unified framework for the joint generation of singing vocals and human motions from text with a multi-modal transformer in an autoregressive fashion.
  • To unify representations of different modalities, we employ a vocal-to-unit model to obtain quantized audio tokens and utilize compositional VQVAEs to get discrete motion tokens.
  • Experimental results show that our framework rivals the performance of specialized single-modality generation systems, setting new benchmarks for joint generation of vocals and motion.
2 Related Work
2.1 Text to Vocal Generation
Text-to-audio Dataset. There exists several singing vocal datasets, yet they each has constraints. For instance, PopCS liu2022diffsinger and OpenSinger huang2021multi are limited to Chinese, while NUS-48E duan2013nus and NHSS sharma2021nhss have only a few hours of songs. Nevertheless, JVS-MuSiC tamaru2020jvs and NUS-48E duan2013nus offer a few hours of songs from dozens of singers, whereas OpenSinger huang2021multi provides a more extensive collection with tens of hours from a single singer. Notably, our dataset represents the first to be specifically curated for rap songs of multiple singers with 108 hours.
Text-to-vocal Models. Recent advancements in text-to-speech (TTS) models, including WavNet oord2016wavenet , FastSpeech 1 and 2 ren2019fastspeech ; ren2020fastspeech , and EATS donahue2020end , have significantly improved the quality of synthesized speech. However, singing voice synthesis (SVS) presents a greater challenge due to its reliance on additional musical scores and lyrics. Recent generation models liu2022diffsinger ; he2023rmssinger ; he2023rmssinger ; zhang2022wesinger ; kim2023muse perform excellently in generating singing voices. However, their performance tends to deteriorate when encountering out-of-distribution data. To handle this, StyleSinger zhang2023stylesinger introduces a Residual Style Adaptor and an Uncertainty Modeling Layer Normalization to handle this.
2.2 Text to Motion Generation
Text-to-motion Dataset. Current text-motion datasets, such as KIT plappert2016kit , AMASS mahmood2019amass , and HumanML3D guo2022generating , are constrained by limited data coverage, typically spanning only tens of hours and lacking whole-body motion representation. To overcome these shortcomings, Motion-X lin2024motion was developed, providing an extensive dataset that encompasses whole-body motions.
Text-to-motion Models. Text-to-motion models petrovich2022temos ; zhang2024motiondiffuse ; guo2022tm2t ; guo2022generating ; lu2023humantomato ; chen2023executing have gained popularity for their user convenience. Recent advancements have focused on diffusion models zhang2024motiondiffuse ; kim2023flame ; tevet2022human , which, unlike deterministic generation models ahuja2019language2pose ; ghosh2021synthesis , enable finer-grained and diverse output generation. Chen et al. chen2023executing proposed a motion latent-based diffusion model to enhance generative quality and reduce computational costs. However, the misalignment between natural language and human motions presents challenges. Lu et al. lu2023humantomato introduced the Text-aligned Whole-body Motion generation framework to address these issues by generating high-quality, diverse, and coherent facial expressions, hand gestures, and body motions simultaneously.
2.3 Audio to Motion Generation
Audio-to-motion Dataset. Co-Speech Datasets can be classified into 1: pseudo-labeled (PGT) and 2: motion-captured (mocap). PGT datasets cao2017realtime ; habibie2021learning ; ginosar2019learning ; yi2023generating utilize disconnected 2D or 3D key points to represent the body, allowing extracting at a lower cost but with limited accuracy. On the other hand, mocap datasets provide annotations for limited body parts, with some focusing solely on the head fanelli20103 ; cudeiro2019capture ; wuu2022multiface or the body takeuchi2017creating ; ferstl2018investigating . Differing from them, BEATX liu2023emage contains mesh data of both head and body. However, these datasets typically focus on speech to motion generation. Our RapVerse, in contrast, contains paired text-vocal-motion data, enabling simultaneous motion and vocal generation.
Audio-to-motion Models. Audio-to-motion models taylor2017deep ; shen2023difftalk ; zhang2023sadtalker ; liu2023emage ; yi2023generating aim to produce human motion from audio innputs. Recognizing the intricate relationship between audio and human face, TalkSHOW yi2023generating separately generates face parts and other body parts. However, this approach exhibits limitations, such as the absence of generating lower body and global motion. EMAGE liu2023emage utilizes masked gestures together with vocals for a generation. Our model, however, goes a step further by being the first to generate paired audio-motion data directly from text.
Dataset Language ##\##Singers Hours
NUS-48E duan2013nus English 12 1.91
NHSS sharma2021nhss English 10 7
JVS-MuSiC tamaru2020jvs Japanese 100 2.28
Tohoku Kiritan ogawa2021tohoku Japanese 1 1
PopCS liu2022diffsinger Chinese 1 5.89
OpenSinger huang2021multi Chinese 66 50
Opencpop wang2022opencpop Chinese 1 5.25
M4Singer zhang2022m4singer Chinese 20 29.77
Ours English 32 108.44
(a) Audio Datasets
Dataset Modality Whole Hours
KIT-ML plappert2016kit Text
11.2
AMASS mahmood2019amass Text
40.0
HumanML3D guo2022generating Text
28.6
Motion-X lin2024motion Text
127.1
HA2G liu2022diffsinger Audio
33
Yoon et.al lu2023co Audio
30
Speech2Gesture ginosar2019learning Audio
144
Talkshow yi2023generating Audio
27.0
Ours Text-Audio
26.8
(b) Motion Datasets
Table 1: Comparison of Audio and Motion Datasets. (a) compares our Rap-Vocal Subset with existing singing vocal datasets. (b) compares our Rap-Motion Subset with existing human motion datasets. : “Whole” means whole-body motion including face, body, and hand.
Figure 2: RapVerse dataset collection pipeline. There are two pathways for rap-vocal and rap-motion subsets, respectively.
3 RapVerse Dataset
In this section, we introduce RapVerse, a large rap music motion dataset containing synchronized singing vocals, textual lyrics and whole-body human motions. A comparison of our dataset with existing datasets is shown at Table. 1. The RapVerse dataset is divided into two subsets to cater to a broad range of research needs: a Rap-Vocal subset and a Rap-Motion subset. The overall collection pipeline of RapVerse is shown at Fig. 2.
3.1 Rap-Vocal Subset
The Rap-Vocal subset contains 108.44 hours of high-quality English singing voice in the rap genre with paired lyrics. We will introduce each step in detail.
Data Crawling. In a bid to obtain a large number of rap songs and corresponding lyrics from the Internet, we utilize Spotdl and Spotipy to collect songs, lyrics, and metadata of different rap singers. To ensure the quality of the dataset, we perform cleaning on the crawled songs by removing songs with misaligned lyrics and filtering out songs that are too long or too short.
Vocal and Background Music Separation. Since the crawled songs are mixed with rapping vocals and background music, and we aim to synthesize singing vocals from separated clean data, we utilize Spleeter hennequin2020spleeter , the state-of-the-art open-source vocal-music source separation tool to separate and extract rap vocal voices and accompanying background music from the collected songs. Following ren2020deepsinger , we normalize the loudness of the vocal voices to a fixed loudness level.
Vocal Data Processing. The raw crawled lyrics from the Internet are in inconsistent formats, we conduct data cleaning on the lyrics by removing meta information (singer, composer, song name, bridging words, and special symbols). To ensure that the lyrics are aligned with the singing vocals, we collect lyrics only with the correct timestamps of each sentence, and we separate each song into around 10-second to 20-second segments for model training.
3.2 Rap-Motion Subset
The Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters pavlakos2019expressive , synchronous singing vocals, and corresponding lyrics. We introduce the collection pipeline of this subset as follows.
Data Crawling. We crawled over 1000 studio performance videos from YouTube under the Common Creative License. We filter out low-quality videos manually to ensure the videos meet the following criteria: stable camera work, performers centered in the frame, clear visibility of the performer’s entire body to capture detailed motion data, and high-quality audio for accurate vocal analysis.
Audio Data Processing. Similar to the Rap-Vocal subset, we leverage Spleeter spleeter2020 to isolate singing vocals from accompanying music. Given that YouTube videos typically lack paired lyrics, we employ an ASR model, Whisper radford2023robust , to accurately transcribe vocals into corresponding text.
Video Data Processing. To ensure the collection of high-quality video clips for motion annotation, we implemented a semi-automatic process to filter out undesirable content, such as advertisements, transition frames, changes in shots, and flashing lights. Initially, we applied YOLO redmon2018yolov3 for human detection to discard frames where no humans were detected. Subsequently, we utilized RAFT teed2020raft to assess the motion magnitude, employing a threshold to eliminate frames affected by camera instability. We then perform meticulous manual curation on the extracted clips, retaining only those of the highest quality. Finally, we follow the pipeline of the optimized-based method Motion-X lin2024motion to extract 3D whole-body meshes from monocular videos. Specifically, we adopt the SMPL-X pavlakos2019expressive for motion representations, given a T-frame video clip, the corresponding pose states $\mathcal{M}$ are represented as:
=$\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{b},\mathcal{M}_{h},\zeta,\epsilon\}$ (1)
where $\mathcal{M}_{f}\in\mathbb{R}^{T\times 3},\mathcal{M}_{b}\in\mathbb{R}^{T\times 63}$ and $\mathcal{M}_{h}\in\mathbb{R}^{T\times 90}$ denotes jaw poses, body poses and hand poses, respectively. $\zeta\in\mathbb{R}^{T\times 100}$ and $\epsilon\in\mathbb{R}^{T\times 3}$ are the facial expression and global translation.
Figure 3: Pipeline overview. We first pre-train all tokenizers on vocal-only and motion-only data. After we have pretrained the modality tokenizers, we can unify text, vocal, and motion in the same token space. We adopt a mixing organizing algorithm for input tokens to align via the temporal axis. These mixed input tokens are fed into the large Text-Motion-Audio foundation model to train on token prediction tasks, guided by the encoded features from textual input.
4 Method
Given a piece of lyrics text, our goal is to generate rap-style vocals and whole-body motions, including body movements, hand gestures, and facial expressions that resonate with the lyrics. With the help of our RapVerse dataset, we propose a novel framework that not only represents texts, vocals, and motions as unified token forms but also integrates token modeling in a unified model. As illustrated in Fig. 3, our model consists of multiple tokenizers for motion (Sec. 4.2) and vocal (Sec. 4.3) token conversions, as well as a general Large Text-Motion-Audio Foundation Model (Sec. 4.4) that targets for audio token synthesize and motion token creation, based on rap lyrics.
4.1 Problem Formulation
Let $\mathcal{L}$ be the rap lyric, such as “I going to make it a flashback set a picture from way back.”, our model will compose the text-related vocal $\mathcal{V}\in\mathbb{R}^{t\times d_{v}}$ and whole-body motion $\mathcal{M}\in\mathbb{R}^{t\times d_{m}}$, where $t$, $d_{v}$, $d_{m}$ denote the temporal length and the feature dimensions for vocal and motion in every time unit, respectively. Modality tokenizer encoders $\phi^{\mathcal{X}}(\mathcal{T}\mid\mathcal{X})$ and decoders $\varphi^{\mathcal{X}}(\mathcal{X}\mid\mathcal{T})$ will bridge each modality with the token format $\mathcal{T}$ utilizing different tokenizers, where $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$. The large foundation model $\psi^{\mathcal{X}}$ will incorporate tokens in unified modeling. The whole problem of coherent vocal and model generation via lyric text can be formulated as follows:
$\Theta^{*}=\mathop{\arg\max}_{\Theta}\mathbb{P}_{\Theta}(\mathcal{V},\mathcal{M}\mid\mathcal{L}),$ (2)
where $\Theta$ denotes the whole model parameters and $\mathbb{P}_{\Theta}(\cdot)$ denotes the model distributions.
4.2 Motion VQ-VAE Tokenizer
Similar to prior arts, a motion VQ-VAE is adopted to convert 3D human motions into discrete tokens, where the tokenizer encoder $\phi$ generates motion tokens enriched with high-frequency information, while the decoder $\varphi$ seamlessly reconstructs discrete tokens into continuous motion sequences. Specifically, the tokenizer will maintain a learnable codebook $\mathcal{C}$, which firstly encodes the input motion features $\mathcal{M}$ to vectors $z^{\mathcal{M}}$, and then the motion quantizer $\mathcal{Q}^{\mathcal{M}}$ will look up the nearest neighbor in codebook $\mathcal{C}^{\mathcal{M}}=\{c_{k}^{\mathcal{M}}\}_{k=1}^{d_{c}}$ for the motion representation. The quantized code can be calculated as
$Q(z^{\mathcal{M}};\mathcal{C}^{\mathcal{M}})=\mathop{\arg\min}_{k}\|z-c_{k}^{\mathcal{M}}\|_{2}.$ (3)
The tokenizer decoder $\varphi$ will project the quantized code back into the human motion after quantization. The reconstruction objective in the tokenizer training stage can be represented as
$\mathbb{E}_{\mathcal{T}\sim\phi_{\mathcal{M}}(\mathcal{T}\mid\mathcal{M})}\left[\log\varphi_{\mathcal{M}}(\mathcal{M}\mid\mathcal{T})\right].$ (4)
As the whole-body motion $\mathcal{M}$ can be decomposed into face $\mathcal{M}_{f}$, body $\mathcal{M}_{b}$, and hand $\mathcal{M}_{h}$, where $\mathcal{M}\in\mathbb{R}^{t\times d},d_{f}+d_{b}+d_{h}=d_{m}$, we build three tokenizers to obtain tokens for these three parts.
4.3 Vocal2unit Audio Tokenizer
Overall, we leverage the self-supervised framework polyak2021speech in speech resynthesis domain to learn vocal representations from the audio sequences. Specifically, we train a Vocal2unit audio tokenizer to build a discrete tokenized representation for the human singing voice. The vocal tokenizer consists of three encoders and a vocoder. The encoders include three different parts: (1) the semantic encoder; (2) the F0 encoder; and (3) the singer encoder. We will introduce each component of the model separately.
Semantic Encoder. In order to extract the semantic information of the audio into discrete tokens, we use a pre-trained Hubert encoder hsu2021hubert as the main component of our semantic encoder network $E^{h}$. Given a raw audio sequence $v$, the output of the Hubert encoder is continuous representation. Then we apply a K-means algorithm on the representations to finally generate discrete units $z^{\mathcal{S}}=\{z^{\mathcal{S}}_{i}\}_{i=1}^{L_{s}}$ where $z^{\mathcal{S}}_{i}\in\{1,2,\cdots,K\}$. $L_{s}$ is the length of hubert token sequence, and $K$ represents the number of clusters in K-means algorithm.
F0 Encoder. We use the YAAPT algorithm kasi2002yet to extract the F0 from the input audio signal $v$, which is then fed into an encoder $E_{p}$ to generated low-frequency F0 pitch representation. The encoder $E^{p}$ is trained using a VQ-VAE framework with Exponential Moving Average updates similar to dhariwal2020jukebox , maintaining a learnable codebook $\mathcal{C}^{p}=\{c_{k}^{p}\}_{k=1}^{d_{c}}$, where $d_{c}$ is the number of codes in the codebook. The raw F0 sequence is first preprocessed according to the singer, converted into a sequence of latent vectors, then mapped to its nearest neighbors in the codebook $c^{p}$, which could be represented using integer number $k\in\{1,2,\cdots,d_{c}\}$ as its index in the codebook. The index sequence will be further formatted as the token <pitch_k> to produce pitch token sequence $z^{\mathcal{P}}=\{z^{\mathcal{P}}_{i}\}_{i=1}^{L_{p}}$, where $L_{p}$ is the length of the pitch token sequence. Finally, the decoder will reconstruct the F0 representation leveraging the pitch token sequence.
Singer Encoder. As the last component of our encoders, the singer encoder $E_{si}$ from heigold2016end ; polyak2021speech is used to extract the Mel-spectrogram from the raw audio sequence and to output a singer representation $z^{\mathcal{I}}\in\mathbb{R}^{256}$. The singer embedding only depends on the singer and is trained globally.
Vocoder. In order to decode the vocal signal from the discrete tokens, we adopt a modified version of the HiFi-GAN neural vocoder similar to the one in polyak2021speech , consisting of a generator and multiple discriminators. The generator takes the encoded discrete representation $z^{\mathcal{S}}$ and $z^{\mathcal{P}}$, and the singer embedding vector $z^{\mathcal{I}}$ as inputs. After a lookup table and a series of blocks composed of transposed convolution and a residual block with dilated layers, it finally outputs waveform audio. The discriminators include multiple sub-discriminators, analyzing on different scales and different periodic structures.
4.4 General Auto-regressive Modeling
Model Architectures. The whole model of our general modeling pipeline consists of several pre-trained vocal and motion tokenizers and a large Text-Motion-Audio foundation model. The vocal and motion tokenizers are based on VQ-VAE architectures, enabling us to represent these two modalities as discrete tokens. We also adopt a T5-Tokenizer raffel2020exploring to convert lyric text into tokens, so that all three modalities $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$ are united in the token space $\mathcal{T}^{\mathcal{X}}=\{t_{k}^{\mathcal{X}}\}_{k=1}^{\mathcal{N}^{\mathcal{X}}}$, where $\mathcal{N}^{\mathcal{X}}$ is the token sequence length for each modality. The large foundation model, a temporal transformer based on the decoder-only architecture, performs the next-token-prediction task based on textual features encoded by the T5-Encoder.
Specifically, each token is generated after the start token by the probability distribution from the foundation model $p_{\delta}(t^{\mathcal{X}};h^{\mathcal{L}})=\prod\limits_{i}p_{\delta}(t_{i}^{\mathcal{X}}\mid t_{<i}^{\mathcal{X}};h_{\mathcal{L}})$. During the training period, we calculate the cross-entropy loss to measure the likelihood between ground truth tokens and predicted token probabilities as
$-\sum\limits_{i=1}^{\mathcal{N^{\mathcal{X}}}}\log p_{\delta}(t^{\mathcal{X}}_{i};h^{\mathcal{L}}).$ (5)
After optimizing via this training objective, our model learns to predict the next token, which can be decoded into different modality features. This process is similar to text word generation in language models, while the “word” in our method such as <face_0123>, does not have explicit semantic information, but can be decoded into continuous modality features.
Multimodality Tokens Organization. Since we have designed a coherent vocal and motion generation pipeline with multiple token types in our method, including hubert and pitch tokens in rap vocals, and face, body, hand in whole-body motions, the organization method of all these tokens in one sentence is important for training. Similar to previous works lu2023humantomato ; jiang2024motiongpt , we adopt an interleaved style within each modality. Specifically, with the input whole-body motion tokens $\mathcal{T}^{\mathcal{M}_{s}}=\{t^{\mathcal{M}_{s}}_{k}\}_{k=1}^{\mathcal{N}^{\mathcal{M}_{s}}}$, where $s=f,b,h$ which denotes face, body and hand respectively, and $\mathcal{N}^{\mathcal{M}_{s}}$ denotes the token length for each part. Therefore the motion token sequence can be formulated as $\mathcal{T}^{\mathcal{M}}=\{t^{\mathcal{M}_{f}}_{1},t^{\mathcal{M}_{b}}_{1},t^{\mathcal{M}_{h}}_{1},t^{\mathcal{M}_{f}}_{2},...\}$. Similarly, the vocal tokens consist of the hubert tokens $t^{\mathcal{V}_{h}}$ and pitch tokens $t^{\mathcal{V}_{p}}$, which can be formulated as $\mathcal{T}^{\mathcal{V}}=\{t^{\mathcal{V}_{h}}_{1},t^{\mathcal{V}_{p}}_{1},t^{\mathcal{V}_{h}}_{2},t^{\mathcal{V}_{p}}_{2},...\}$. This interleaved way of organization allows us to align the timesteps of each part. When performing coherent vocal and motion generation, we mix these two modality sequences sequentially with modality-wise start token as $\mathcal{T}=\{t_{start}^{V},\mathcal{T}^{V},t_{start}^{M},\mathcal{T}^{M}\}$. Note that we put vocal sequences before motion sequences because 1) the vocals are more related to the lyrics, and 2) motion tokens have a relationship with previously generated vocals, e.g. face motions with lips movement are directly related to vocals. Joint training of mixed modalities allows the transformer to capture the temporal dependencies and dynamics inherent in vocal and motion data.
Inference and Decoupling. In the inference stage, we use different start tokens to specify which modality to generate. The textual input is encoded as features to guide token inference. We also adopt a top-k algorithm to control the diversity of the generated content by adjusting the temperature, as generating vocals and motions based on lyrics is a creation process with multiple possible answers. After token prediction, a decoupling algorithm is used to process output tokens to make sure tokens from different modalities are separated and temporally aligned. These discrete tokens will be further decoded into text-aligned vocals and motions.
5 Experiments
In this section, we evaluate our proposed model on our proposed benchmark designed for joint vocal and whole-body motion generation from textual inputs.
5.1 Experimental Setup
Metrics. To evaluate the generation quality of singing vocals, we utilize the Mean Opinion Score (MOS) to gauge the naturalness of the synthesized vocal. For motion synthesis, we evaluate the generation quality of the body hand gestures and the realism of the face, respectively. Specifically, for gesture generation, we use Frechet Inception Distance (FID) based on a feature extractor from guo2022generating to evaluate the distance of feature distributions between the generated and real motions, and Diversity (DIV) metric to assess the motions diversity. For face generation, we compare the vertex MSE xing2023codetalker and the vertex L1 difference LVD yi2023generating . Finally, we adopt Beat Constancy (BC) li2021ai to measure the synchrony of generated motion and singing vocals.
Baselines. We compare the vocal generation quality with the state-of-the-art vocal generation method DiffSinger liu2022diffsinger . And we also adapt the text-to-speech model FastSpeech2 ren2020fastspeech for vocal generation. For motion generation, we compare our method with both text-to-motion methods and audio-to-motion methods. For text-to-motion methods, since there is no existing open-sourced work for text to whole-body motion generation, we compare with transformer-based T2M-GPT zhang2023t2m and MLD chen2023executing for body generation. For the audio-to-motion generation, we compare with Habibie et al. habibie2021learning and the SOTA model Talkshow yi2023generating . We report all the results on RapVerse with an 85%/7.5%/7.5% train/val/test split.
[…]
Text-to-Motion
T2M-GPT zhang2023t2m 18.42 11.75 - - -
MLD chen2023executing 24.52 12.15 - - -
Audio-to-Motion
Habibie et al. habibie2021learning 32.14 10.08 0.476 2.11 9.54
Talkshow yi2023generating 21.10 13.14 0.487 1.93 9.20
Cascaded Result 23.42 12.87 0.479 2.05 9.38
Text-to-Audio+Motion
Ours 18.65 12.19 0.481 2.03 7.23
(a) Motion Generation
[…]
GT 4.45 ± 0.06
Reconstruction 4.02 ± 0.08
FastSpeech2 ren2020fastspeech 3.41 ± 0.18
DiffSinger liu2022diffsinger 3.72 ± 0.12
Ours 3.64 ± 0.15
(b) Vocal Generation
Table 2: Quantitative results of generated motion and vocal. (a) compares with different motion generation baselines, best result is shown in bold and the second-best result is shown in underline. (b) compares with text-to-vocal generation baselines. : The MOS is calculated with 95% confidence intervals of song samples.
5.2 Main Results Analysis
Evaluations on joint vocal and whole-body motion generations. We compared both text-driven and audio-driven motion generation baselines in Table. 2 (a). To be noted, our setting is different from all existing methods in the following ways. First, we use rap lyrics as our textual input instead of motion textual descriptions, which contain direct action prompt words, such as walk and jump; Second we use text to jointly generate both audio and motion, instead of using audio to generate motion as audio-driven methods did. As is demonstrated, our model rivals with both text-to-motion and audio-to-motion methods in terms of metrics measuring body motion quality and face motion accuracy.
Furthermore, the cornerstone of our approach lies in the simultaneous generation of vocals and motion, aiming to achieve temporal alignment between these modalities. This objective is substantiated by our competitive results on the BC metric, which assesses the synchrony between singing vocals and corresponding motions, underscoring our success in closely synchronizing the generation of these two modalities. For the cascaded system, we integrate the text-to-vocal model DiffSinger with the audio-to-motion model Talkshow. Compared with the cascaded system, our joint-generation pipeline demonstrates superior outcomes while also reducing computational demands during both training and inference phases. In the cascaded architectures, errors tend to accumulate through each stage. Specifically, if the text-to-vocal module produces unclear vocals, it subsequently hampers the audio-to-motion model’s ability to generate accurate facial expressions that align with the vocal content.
Evaluations on vocal generations. We have also compared our method against other state-of-the-art text-to-vocal generation baselines in Table. 2 (b). While our unified model is trained to simultaneously generate vocals and motion, a task considerably more complex than generating vocals alone, its vocal generation component still manages to achieve results comparable to those systems designed solely for vocal generations.
[…]
GT 0 12.53 0.499 0 0
Pretrained LLM 48.25 10.05 0.462 3.28 12.15
Single motion token 19.15 11.29 0.477 2.18 10.12
Ours 18.65 12.19 0.481 2.03 7.23
Table 3: Ablation study. We compare with common designs in motion generation frameworks.
5.3 Ablation Study
We present the outcomes of our ablation study in Table. 3. Initially, we explored the integration of a pre-trained large language model raffel2020exploring for multi-modality generation, akin to the approach in  jiang2024motiongpt . However, the efficacy of utilizing pre-trained language models significantly lags behind our tailored design, underscoring that pre-training primarily on linguistic tokens does not facilitate effective prediction across multiple modalities like vocal and motion. Additionally, we study the impact of our compositional VQ-VAEs on motion generation. In contrast, a baseline employing a single VQVAE for the joint quantization of facial, body, and hand movements was implemented. This approach led to a noticeable degradation in performance, particularly marked by a -2.89 decrease in LVD. This decline can be attributed to the preponderance of facial movements in a singer’s performance. Using a single VQ-VAE model for full-body dynamics compromises the detailed representation of facial expressions, which are crucial for realistic and coherent motion synthesis.
6 Conclusion
In this work, we present a new framework for the simultaneous generation of 3D whole-body motions and singing vocals, directly from textual lyrics. To address this challenging task, we first collect RapVerse, a large dataset containing synchronous rap vocals, alongside lyrics and 3D whole-body motions. Utilizing RapVerse, we demonstrate that simply scaling autoregressive transformers across language, audio, and motion yields a coherent generation of singing vocals and 3D holistic human motions. We anticipate that this work will inspire novel avenues in the joint modeling of text, audio, and motion.
Limitations and Future Works. One limitation of RapVerse is that it is currently constrained to the rap music generation, excluding other music genres. However, we want to emphasize that our method is a general framework that could be flexibly utilized in other scenarios for joint audio and motion generation if provided with the datasets. Another important future direction resides in multi-performer audio and motion generation, which could be utilized in contemporary music performances such as virtual live bands.
Appendix A Appendix
The supplementary material is organized as follows: Sec. A.1 provides a website demo to show additional qualitative results; Sec. A.2 presents additional details of the network architectures; Sec. A.3 introduces evaluation metrics; Sec. A.4 presents additional ablation studies; Sec. A.5 shows additional qualitative results; Sec. A.6 discusses broader societal impacts of the work.
A.1 Website Demo
In order to provide more vivid and clear qualitative results, we make a supplemental website demo to demonstrate the generation quality of our proposed system. We encourage readers to view the results at https://vis-www.cs.umass.edu/RapVerse/.
A.2 Implementation Details
Motion Tokenizer. We train three separate Vector Quantized Variational Autoencoders (VQ-VAE) for face, body and hand, respectively. We adopt the same VQ-VAE architecture based on guo2022tm2t ; zhang2023t2m ; jiang2024motiongpt . For loss functions optimizing the motion tokenizers, we use the L1 smooth reconstruction loss, the embedding loss, and the commitment loss. The commitment loss weight is set to 0.02. In line with guo2022tm2t ; jiang2024motiongpt , strategies such as the exponential moving average and the codebook reset technique razavi2019generating are implemented to optimize codebook efficacy throughout the training process. We take 512 for the codebook size and set the dimension of each code to 512. We set the temporal down-sampling rate to 4. We train the VQ-VAEs with a batch size of 256 and a sequence window length of 72. We adopt Adam with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of 0.0002 as the optimizer.
Vocal Tokenizer. For the semantic encoder, we adopt a BASE 12 transformer of HuBERT hsu2021hubert pre-trained on the 969-hour LibriSpeech corpus panayotov2015librispeech . Following lakhotia2021generative ; polyak2021speech , we derive the feature activations from its sixth layer. This process allows the HuBERT model to transform input audio into a 768-dimensional vector space. Subsequently, we employ the k-means algorithm with 500 centroids to get quantized discrete content codes. For the F0 encoder, a VQ-VAE framework is utilized to discretize the F0 signal into quantized F0 tokens. We adopt the Exponential Moving Average updates during training the VQ-VAE following dhariwal2020jukebox ; polyak2021speech . We set the codebook size of the VQ-VAE to 20 entries. Moreover, as the original work directly normalizes the extracted F0 values for each singer respectively, we don’t explicitly use the singer’s statistics but adopt a windowed convolutional layer with both the audio input (sliced into the window size) and singer’s embedding as input. Finally, we adopt a similar architecture as heigold2016end for the singer encoder.
General Auto-regressive Model. The auto-regressive model consists of a T5 Embedder and a Foundation Model. We use a T5-Large Encoder as our Embedder, with 24 layers and 16 heads. The Embedder is freezed during foundation model training. The foundation model is based on the Decoder-only transformer architecture, which has 12 layers and 8 heads. We use Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of 0.0002. We do not use dropout in our training. Our training batch size is 384 for 100 epochs.
A.3 Evaluation Metrics
To evaluate the motion generation quality, we utilize the following metrics:
  1. 1.
    Frechet Inception Distance (FID): This metric measures the distribution discrepancy between the ground truth and generated motions of body and hand gestures. Specifically, we train an autoencoder based on guo2022tm2t as the motion feature extractor.
  2. 2.
    Diversity (DIV): DIV evaluates the diversity of generated motions, where we calculate the variance from the extracted motion features.
  3. 3.
    Vertex MSE: Following xing2023codetalker , we compute the mean L2 error of lip vertices between generated face motions and ground truth.
  4. 4.
    Landmark Velocity Difference (LVD): Introduced by yi2023generating , LVD calculates the velocity difference of generated facial landmarks and ground truth.
  5. 5.
    Beat Constancy (BC) li2021ai : BC evaluates the synchrony of generated motions and singing vocals by calculating the similarity between the rhythm of gestures and audio beat. Specifically, we extract audio beats using librosa mcfee2015librosa , and we compute the kinematic beats as the local minima of the velocity of joints. Then the alignment score is derived from the mean proximity of motion beats to the nearest audio beat.
For the evaluation of singing vocal generation quality, the Mean Opinion Score (MOS) is employed. It reflects the perceived naturalness of the synthesized vocal tones, with human evaluators rating each sample on a scale from 1 to 5, thereby offering a subjective measure of vocal synthesis fidelity.
A.4 Additional Ablation Studies
Ablation on Motion Tokenizer. We study different designs of our motion tokenizer by comparing the reconstruction results. Specifically, we explore VQ-VAEs with different codebook sizes, and study the effect of using a single VQ-VAE to model full-body motions instead of multiple VQ-VAEs for different body parts. As is demonstrated in Table. 4, we find that using separate VQ-VAEs for face, body and hands has lower reconstruction error. And we select a codebook size of 512 for our final model.
[…]
K = 256 63.4 39.1 9.14
K = 512 60.9 35.4 8.90
K = 1024 62.2 36.8 9.08
Single 65.2 38.8 9.32
Ours 60.9 35.4 8.90
Table 4: Evaluation of our motion tokenizer. We follow lu2023humantomato to evaluate the motion reconstruction errors of our motion VQVAE model $V_{m}$. MPJPE and PAMPJPE are measured in millimeters. ACCL indicates acceleration error.
Ablation on Vocal Tokenizer. We also study different designs for our audio tokenizer by comparing the reconstruction results. Specifically, we explore different codebook sizes for the semantic encoder by changing the K-Means number. We also compare the effect with our singer embedding in F0 value postprocessing. We use the following metrics to measure the reconstruction quality of the vocal tokenizer:
  1. 1.
    Character Error Rate (CER): We use Whisper radford2022whisper to transcribe the ground truth and synthesized audios, and then take the corresponding ground truth lyrics as the reference to calculate the CER of the synthesized audios.
  2. 2.
    Gross Pitch Error (GPE): The percentage of pitch estimates that have a deviation of more than 20% around the true pitch. Only frames considered pitched by both the ground truth and the synthesized audio are considered.
  3. 3.
    Voicing Decision Error (VDE) NAKATANI2008203 : The portion of frames with voicing decision error, that is, the results using ground truth and synthesized audio to determine whether the frame is voiced are different.
[…]
GT 41.88 - -
K = 100 76.39 2.64 11.98
K = 500 (Ours) 69.21 2.29 8.84
K = 2000 67.46 2.54 9.08
Original F0 post-process 68.93 2.71 9.19
Table 5: Evaluation of our unit2wav model. We follow polyak2021speech to evaluate the speech resynthesis errors of our unit2wav model. WER, GPE, and VDE, expressed as percentages, indicate the character error rate, the grand pitch error and the voicing decision error.
The analysis of the results indicates that even the original audio exhibits a high CER, which could be attributed to the rapid speech rate associated with rapping. In some instances, the lyrics may not be distinctly recognizable even by human listeners. Upon comparing different codebook sizes, it is observed that they achieve comparable GPE values. This similarity in GPE is expected since the same F0 model is employed across all codebook sizes. The CER, which serves as a direct measure of the semantic information preserved in the code, suggests that larger codebooks tend to retain more semantic information. However, the difference in CER between codebook sizes of $K=500$ and $K=2000$ is minimal. Given that $K=500$ demonstrates better GPE and VDE, we select $K=500$.
Additionally, we ablate a design without the singer embedding in the F0 preprocessing, instead normalizing the F0 values for each singer. It shows that this approach resulted in significantly inferior performance, particularly in pitch prediction, compared to the modified version that includes the singer embedding.
A.5 Additional Qualitative Results
We show additional qualitative results in Fig. 4. Our model adeptly generates comprehensive whole-body motions that embody the essence of the input lyrics. These include authentic gesture movements that resonate with the song’s rhythm and synchronized lip motions that articulate the lyrics.
Figure 4: Additional qualitative results. Our method can generate diverse whole-body motions from input lyrics.
A.6 Broader Impacts
This research contributes to advancements in generating synchronized vocals and human motion from textual lyrics, aiming to enhance virtual agents’ ability to provide immersive and interactive experiences in digital media. The potential positive impact of this work lies in its ability to create more lifelike and engaging virtual performances, such as in virtual concerts and gaming, where characters can perform and react in ways that are deeply resonant with human expressions. This could significantly enhance user engagement in virtual reality settings, and provide innovative solutions in entertainment industries.
However, this capability carries inherent risks of misuse. The technology’s ability to generate realistic human-like actions and singing vocals from mere text raises concerns about its potential to create misleading or deceptive content. For example, this could be exploited to produce fake videos or deepfakes where individuals appear to sing and perform that never actually occurred, which could be used to spread misinformation or harm reputations. Recognizing these risks, it is crucial to advocate for ethical guidelines and robust frameworks to ensure the responsible use of such technologies.

### v3

@@
-RapVerse: Coherent Vocals and Whole-Body Motions Generations from Text
+RapVerse: Coherent Vocals and Whole-Body Motion Generation from Text
 Abstract
@@
 Extensive experiments demonstrate that our unified generation framework not only produces coherent and realistic singing vocals alongside human motions directly from textual inputs, but also rivals the performance of specialized single-modality generation systems, establishing new benchmarks for joint vocal-motion generation.
-The project page is available for research purposes at https://vis-www.cs.umass.edu/RapVerse.
-1 Introduction
+Video demonstration and more information can be found on the project page [^1] https://jiabenchen.github.io/RapVerse/..
 Figure 1: RapVerse.
@@
 Note that the corresponding video frames are just shown for reference.
-In the evolving landscape of multi-modal content generation in terms of sound and motion, significant strides have been made in individual modalities, including text-to-music schneider2023mo ; agostinelli2023musiclm ; huang2023noise2music , text-to-vocal liu2022diffsinger , text-to-motion guo2022generating ; zhang2023t2m ; chen2023executing ; jiang2024motiongpt ; lu2023humantomato , and audio-to-motion yi2023generating ; habibie2021learning ; liu2023emage generation.
+1 Introduction
+In the evolving landscape of multi-modal content generation in terms of sound and motion, significant strides have been made in individual modalities, including text-to-music [schneider2023mo, agostinelli2023musiclm, huang2023noise2music], text-to-vocal [liu2022diffsinger], text-to-motion [guo2022generating, zhang2023t2m, chen2023executing, jiang2024motiongpt, lu2023humantomato], and audio-to-motion [yi2023generating, habibie2021learning, liu2023emage] generation.
 These developments have paved the way for creating more dynamic and interactive digital content.
 Despite these advancements, existing works predominantly operate in silos, addressing each modality in isolation.
-However, there’s strong psychological evidence that for human beings, the generation of sound and motion are highly related and coupled lakoff1980metaphors .
+However, there’s strong psychological evidence that for human beings, the generation of sound and motion are highly related and coupled [lakoff1980metaphors].
 A unified system for joint generation allows for a more expressive and nuanced communication of emotions, intentions, and context, where the generation of one modality could guide and assist the other in a coherent and efficient way.
@@
 Firstly, we consider textual lyrics as the proper form of inputs for the unified system, since text provides a highly expressive, interpretable flexible means of conveying information by human beings, and could serve as a bridge between various modalities.
-Previous efforts explore scores liu2022diffsinger , action commands zhang2023t2m ; chen2023executing ; jiang2024motiongpt , or audio signals yi2023generating as inputs, which are inferior to textual inputs in terms of semantic richness, expressiveness and flexible integration of different modalities.
+Previous efforts explore scores [liu2022diffsinger], action commands [zhang2023t2m, chen2023executing, jiang2024motiongpt], or audio signals [yi2023generating] as inputs, which are inferior to textual inputs in terms of semantic richness, expressiveness and flexible integration of different modalities.
 Secondly, we reckon that a joint generation system that could produce multi-modal outputs simultaneously is better than a cascaded system that executes the single-modal generation sequentially.
@@
 In response to these challenges, we have curated RapVerse, a large-scale dataset featuring a comprehensive collection of lyrics, singing vocals, and 3D whole-body motions.
-Despite the existence of datasets available for text-to-vocal liu2022diffsinger ; huang2021multi ; duan2013nus ; sharma2021nhss , text-to-motion plappert2016kit ; mahmood2019amass ; guo2022generating ; lin2024motion , and audio-to-motion cao2017realtime ; habibie2021learning ; ginosar2019learning ; fanelli20103 ; cudeiro2019capture ; wuu2022multiface , the landscape lacks a unified dataset that encapsulates singing vocals, whole-body motion, and lyrics simultaneously.
-Most notably, large text-to-vocal datasets huang2021multi ; zhang2022m4singer are predominantly in Chinese, limiting their applicability for English language research and lacking any motion data.
-And text-to-motion datasets plappert2016kit ; guo2022generating ; lin2024motion typically focus on text descriptions of specific actions paired with corresponding motions without audio data, often not covering whole body movements.
-Moreover, audio-to-motion datasets liu2022diffsinger ; lu2023co focus primarily on speech rather than singing.
+Despite the existence of datasets available for text-to-vocal [liu2022diffsinger, huang2021multi, duan2013nus, sharma2021nhss], text-to-motion [plappert2016kit, mahmood2019amass, guo2022generating, lin2024motion], and audio-to-motion [cao2017realtime, habibie2021learning, ginosar2019learning, fanelli20103, cudeiro2019capture, wuu2022multiface], the landscape lacks a unified dataset that encapsulates singing vocals, whole-body motion, and lyrics simultaneously.
+Most notably, large text-to-vocal datasets [huang2021multi, zhang2022m4singer] are predominantly in Chinese, limiting their applicability for English language research and lacking any motion data.
+And text-to-motion datasets [plappert2016kit, guo2022generating, lin2024motion] typically focus on text descriptions of specific actions paired with corresponding motions without audio data, often not covering whole body movements.
+Moreover, audio-to-motion datasets [liu2022diffsinger, lu2023co] focus primarily on speech rather than singing.
 A comparison of existing related datasets is demonstrated in Table.
@@
 Paired lyrics and vocals are crawled from the Internet from 32 singers, with careful cleaning and post-processing.
-On the other hand, the Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters pavlakos2019expressive using the annotation pipeline of Motion-X lin2024motion , synchronous singing vocals and corresponding lyrics.
+On the other hand, the Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters [pavlakos2019expressive] using the annotation pipeline of Motion-X [lin2024motion], synchronous singing vocals and corresponding lyrics.
 With the RapVerse dataset, we explore how far we can push by simply scaling autoregressive multimodal transformers with language, audio, and motion for a coherent and realistic generation of vocals and whole-body human motions.
 To this end, we unify different modalities as token representations.
-Specifically, three VQVAEs van2017neural are utilized to compress whole-body motion sequences into three-level discrete tokens for head, body, and hand, respectively.
-For vocal generation, previous works min2021meta ; donahue2020end ; liu2022diffsinger ; min2021meta share a common paradigm, producing mel-spectrograms of audio signals from input textual features and additional music score information, following with a vocoder oord2016wavenet ; valin2019lpcnet ; yang2021multi to reconstruct the phase.
-We draw inspiration from the speech resynthesis domain polyak2021speech , and learn a self-supervised discrete representation to quantize raw audio signal into discrete tokens while preserving the vocal content and prosodic information.
+Specifically, three VQVAEs [van2017neural] are utilized to compress whole-body motion sequences into three-level discrete tokens for head, body, and hand, respectively.
+For vocal generation, previous works [min2021meta, donahue2020end, liu2022diffsinger, min2021meta] share a common paradigm, producing mel-spectrograms of audio signals from input textual features and additional music score information, following with a vocoder [oord2016wavenet, valin2019lpcnet, yang2021multi] to reconstruct the phase.
+We draw inspiration from the speech resynthesis domain [polyak2021speech], and learn a self-supervised discrete representation to quantize raw audio signal into discrete tokens while preserving the vocal content and prosodic information.
 Then, with all the inputs in discrete representations, we leverage a transformer to predict the discrete codes of audio and motion in an autoregressive fashion.
@@
 To sum up, this paper makes the following contributions:
-  • We release RapVerse, a large dataset featuring synchronous singing vocals, lyrics, and high-quality 3D holistic SMPL-X parameters.
+  • We collect RapVerse, a large dataset featuring synchronous singing vocals, lyrics, and high-quality 3D holistic SMPL-X parameters.
   • We design a simple but effective unified framework for the joint generation of singing vocals and human motions from text with a multi-modal transformer in an autoregressive fashion.
@@
 There exists several singing vocal datasets, yet they each has constraints.
-For instance, PopCS liu2022diffsinger and OpenSinger huang2021multi are limited to Chinese, while NUS-48E duan2013nus and NHSS sharma2021nhss have only a few hours of songs.
-Nevertheless, JVS-MuSiC tamaru2020jvs and NUS-48E duan2013nus offer a few hours of songs from dozens of singers, whereas OpenSinger huang2021multi provides a more extensive collection with tens of hours from a single singer.
+For instance, PopCS [liu2022diffsinger] and OpenSinger [huang2021multi] are limited to Chinese, while NUS-48E [duan2013nus] and NHSS [sharma2021nhss] have only a few hours of songs.
+Nevertheless, JVS-MuSiC [tamaru2020jvs] and NUS-48E [duan2013nus] offer a few hours of songs from dozens of singers, whereas OpenSinger [huang2021multi] provides a more extensive collection with tens of hours from a single singer.
 Notably, our dataset represents the first to be specifically curated for rap songs of multiple singers with 108 hours.
 Text-to-vocal Models.
-Recent advancements in text-to-speech (TTS) models, including WavNet oord2016wavenet , FastSpeech 1 and 2 ren2019fastspeech ; ren2020fastspeech , and EATS donahue2020end , have significantly improved the quality of synthesized speech.
+Recent advancements in text-to-speech (TTS) models, including WavNet [oord2016wavenet], FastSpeech 1 and 2 [ren2019fastspeech, ren2020fastspeech], and EATS [donahue2020end], have significantly improved the quality of synthesized speech.
 However, singing voice synthesis (SVS) presents a greater challenge due to its reliance on additional musical scores and lyrics.
-Recent generation models liu2022diffsinger ; he2023rmssinger ; he2023rmssinger ; zhang2022wesinger ; kim2023muse perform excellently in generating singing voices.
+Recent generation models [liu2022diffsinger, he2023rmssinger, he2023rmssinger, zhang2022wesinger, kim2023muse] perform excellently in generating singing voices.
 However, their performance tends to deteriorate when encountering out-of-distribution data.
-To handle this, StyleSinger zhang2023stylesinger introduces a Residual Style Adaptor and an Uncertainty Modeling Layer Normalization to handle this.
+To handle this, StyleSinger [zhang2023stylesinger] introduces a Residual Style Adaptor and an Uncertainty Modeling Layer Normalization to handle this.
+Dataset Language #Singers Hours
+NUS-48E [duan2013nus] English 12 1.91
+NHSS [sharma2021nhss] English 10 7
+JVS-MuSiC [tamaru2020jvs] Japanese 100 2.28
+Tohoku Kiritan [ogawa2021tohoku] Japanese 1 1
+PopCS [liu2022diffsinger] Chinese 1 5.89
+OpenSinger [huang2021multi] Chinese 66 50
+Opencpop [wang2022opencpop] Chinese 1 5.25
+M4Singer [zhang2022m4singer] Chinese 20 29.77
+Ours English 32 108.44
+Dataset Modality Whole Hours
+KIT-ML [plappert2016kit] Text
+11.2
+AMASS [mahmood2019amass] Text
+40.0
+HumanML3D [guo2022generating] Text
+28.6
+Motion-X [lin2024motion] Text
+127.1
+HA2G [liu2022diffsinger] Audio
+33
+Yoon et.al [lu2023co] Audio
+30
+Speech2Gesture [ginosar2019learning] Audio
+144
+Talkshow [yi2023generating] Audio
+27.0
+Ours Text-Audio
+26.8
+Table 1: Comparison of Audio and Motion Datasets.
+(a) compares our Rap-Vocal Subset with existing singing vocal datasets.
+(b) compares our Rap-Motion Subset with existing motion datasets.
+: “Whole” means whole-body motion including face, body, and hand.
+Figure 2: RapVerse dataset collection pipeline. There are two pathways for rap-vocal and rap-motion subsets, respectively.
 2.2 Text to Motion Generation
 Text-to-motion Dataset.
-Current text-motion datasets, such as KIT plappert2016kit , AMASS mahmood2019amass , and HumanML3D guo2022generating , are constrained by limited data coverage, typically spanning only tens of hours and lacking whole-body motion representation.
-To overcome these shortcomings, Motion-X lin2024motion was developed, providing an extensive dataset that encompasses whole-body motions.
+Current text-motion datasets, such as KIT [plappert2016kit], AMASS [mahmood2019amass], and HumanML3D [guo2022generating], are constrained by limited data coverage, typically spanning only tens of hours and lacking whole-body motion representation.
+To overcome these shortcomings, Motion-X [lin2024motion] was developed, providing an extensive dataset that encompasses whole-body motions.
 Text-to-motion Models.
-Text-to-motion models petrovich2022temos ; zhang2024motiondiffuse ; guo2022tm2t ; guo2022generating ; lu2023humantomato ; chen2023executing have gained popularity for their user convenience.
-Recent advancements have focused on diffusion models zhang2024motiondiffuse ; kim2023flame ; tevet2022human , which, unlike deterministic generation models ahuja2019language2pose ; ghosh2021synthesis , enable finer-grained and diverse output generation.
+Text-to-motion models [petrovich2022temos, zhang2024motiondiffuse, guo2022tm2t, guo2022generating, lu2023humantomato, chen2023executing] have gained popularity for their user convenience.
+Recent advancements have focused on diffusion models [zhang2024motiondiffuse, kim2023flame, tevet2022human], which, unlike deterministic generation models [ahuja2019language2pose, ghosh2021synthesis], enable finer-grained and diverse output generation.
 Chen et al.
-chen2023executing proposed a motion latent-based diffusion model to enhance generative quality and reduce computational costs.
+[chen2023executing] proposed a motion latent-based diffusion model to enhance generative quality and reduce computational costs.
 However, the misalignment between natural language and human motions presents challenges.
 Lu et al.
-lu2023humantomato introduced the Text-aligned Whole-body Motion generation framework to address these issues by generating high-quality, diverse, and coherent facial expressions, hand gestures, and body motions simultaneously.
+[lu2023humantomato] introduced the Text-aligned Whole-body Motion generation framework to address these issues by generating high-quality, diverse, and coherent facial expressions, hand gestures, and body motions simultaneously.
 2.3 Audio to Motion Generation
@@
 Co-Speech Datasets can be classified into 1: pseudo-labeled (PGT) and 2: motion-captured (mocap).
-PGT datasets cao2017realtime ; habibie2021learning ; ginosar2019learning ; yi2023generating utilize disconnected 2D or 3D key points to represent the body, allowing extracting at a lower cost but with limited accuracy.
-On the other hand, mocap datasets provide annotations for limited body parts, with some focusing solely on the head fanelli20103 ; cudeiro2019capture ; wuu2022multiface or the body takeuchi2017creating ; ferstl2018investigating .
-Differing from them, BEATX liu2023emage contains mesh data of both head and body.
+PGT datasets [cao2017realtime, habibie2021learning, ginosar2019learning, yi2023generating] utilize disconnected 2D or 3D key points to represent the body, allowing extracting at a lower cost but with limited accuracy.
+On the other hand, mocap datasets provide annotations for limited body parts, with some focusing solely on the head [fanelli20103, cudeiro2019capture, wuu2022multiface] or the body [takeuchi2017creating, ferstl2018investigating].
+Differing from them, BEATX [liu2023emage] contains mesh data of both head and body.
 However, these datasets typically focus on speech to motion generation.
@@
 Audio-to-motion Models.
-Audio-to-motion models taylor2017deep ; shen2023difftalk ; zhang2023sadtalker ; liu2023emage ; yi2023generating aim to produce human motion from audio innputs.
-Recognizing the intricate relationship between audio and human face, TalkSHOW yi2023generating separately generates face parts and other body parts.
-However, this approach exhibits limitations, such as the absence of generating lower body and global motion.
-EMAGE liu2023emage utilizes masked gestures together with vocals for a generation.
+Audio-to-motion models [taylor2017deep, shen2023difftalk, zhang2023sadtalker, liu2023emage, yi2023generating, alexanderson2023listen, bian2024adding, chen2024diffsheg] aim to produce human motion from audio innputs.
+Recognizing the intricate relationship between audio and human face, TalkSHOW [yi2023generating] separately generates face parts and other body parts.
+However, this approach exhibits limitations, such as the absence of generating lower body motion.
+Motioncraft [bian2024adding] proposes a DiT structure for motion generation.
+EMAGE [liu2023emage] utilizes masked gestures together with vocals for a generation.
 Our model, however, goes a step further by being the first to generate paired audio-motion data directly from text.
-Dataset Language ##\##Singers Hours
-NUS-48E duan2013nus English 12 1.91
-NHSS sharma2021nhss English 10 7
-JVS-MuSiC tamaru2020jvs Japanese 100 2.28
-Tohoku Kiritan ogawa2021tohoku Japanese 1 1
-PopCS liu2022diffsinger Chinese 1 5.89
-OpenSinger huang2021multi Chinese 66 50
-Opencpop wang2022opencpop Chinese 1 5.25
-M4Singer zhang2022m4singer Chinese 20 29.77
-Ours English 32 108.44
-(a) Audio Datasets
-Dataset Modality Whole Hours
-KIT-ML plappert2016kit Text
-11.2
-AMASS mahmood2019amass Text
-40.0
-HumanML3D guo2022generating Text
-28.6
-Motion-X lin2024motion Text
-127.1
-HA2G liu2022diffsinger Audio
-33
-Yoon et.al lu2023co Audio
-30
-Speech2Gesture ginosar2019learning Audio
-144
-Talkshow yi2023generating Audio
-27.0
-Ours Text-Audio
-26.8
-(b) Motion Datasets
-Table 1: Comparison of Audio and Motion Datasets.
-(a) compares our Rap-Vocal Subset with existing singing vocal datasets.
-(b) compares our Rap-Motion Subset with existing human motion datasets.
-: “Whole” means whole-body motion including face, body, and hand.
-Figure 2: RapVerse dataset collection pipeline. There are two pathways for rap-vocal and rap-motion subsets, respectively.
 3 RapVerse Dataset
@@
 Vocal and Background Music Separation.
-Since the crawled songs are mixed with rapping vocals and background music, and we aim to synthesize singing vocals from separated clean data, we utilize Spleeter hennequin2020spleeter , the state-of-the-art open-source vocal-music source separation tool to separate and extract rap vocal voices and accompanying background music from the collected songs.
-Following ren2020deepsinger , we normalize the loudness of the vocal voices to a fixed loudness level.
+Since the crawled songs are mixed with rapping vocals and background music, and we aim to synthesize singing vocals from separated clean data, we utilize Spleeter [hennequin2020spleeter], the state-of-the-art open-source vocal-music source separation tool to separate and extract rap vocal voices and accompanying background music from the collected songs.
+Following [ren2020deepsinger], we normalize the loudness of the vocal voices to a fixed loudness level.
 Vocal Data Processing.
@@
 3.2 Rap-Motion Subset
-The Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters pavlakos2019expressive , synchronous singing vocals, and corresponding lyrics.
+The Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters [pavlakos2019expressive], synchronous singing vocals, and corresponding lyrics.
 We introduce the collection pipeline of this subset as follows.
@@
 Audio Data Processing.
-Similar to the Rap-Vocal subset, we leverage Spleeter spleeter2020 to isolate singing vocals from accompanying music.
-Given that YouTube videos typically lack paired lyrics, we employ an ASR model, Whisper radford2023robust , to accurately transcribe vocals into corresponding text.
+Similar to the Rap-Vocal subset, we leverage Spleeter [spleeter2020] to isolate singing vocals from accompanying music.
+Given that YouTube videos typically lack paired lyrics, we employ an ASR model, Whisper [radford2023robust], to accurately transcribe vocals into corresponding text.
 Video Data Processing.
 To ensure the collection of high-quality video clips for motion annotation, we implemented a semi-automatic process to filter out undesirable content, such as advertisements, transition frames, changes in shots, and flashing lights.
-Initially, we applied YOLO redmon2018yolov3 for human detection to discard frames where no humans were detected.
-Subsequently, we utilized RAFT teed2020raft to assess the motion magnitude, employing a threshold to eliminate frames affected by camera instability.
+Initially, we applied YOLO [redmon2018yolov3] for human detection to discard frames where no humans were detected.
+Subsequently, we utilized RAFT [teed2020raft] to assess the motion magnitude, employing a threshold to eliminate frames affected by camera instability.
 We then perform meticulous manual curation on the extracted clips, retaining only those of the highest quality.
-Finally, we follow the pipeline of the optimized-based method Motion-X lin2024motion to extract 3D whole-body meshes from monocular videos.
-Specifically, we adopt the SMPL-X pavlakos2019expressive for motion representations, given a T-frame video clip, the corresponding pose states $\mathcal{M}$ are represented as:
-=$\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{b},\mathcal{M}_{h},\zeta,\epsilon\}$ (1)
+Finally, we follow the pipeline of the optimized-based method Motion-X [lin2024motion] to extract 3D whole-body meshes from monocular videos.
+Specifically, we adopt the SMPL-X [pavlakos2019expressive] for motion representations, given a T-frame video clip, the corresponding pose states $\mathcal{M}$ are represented as:
+$\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{b},\mathcal{M}_{h},\zeta,\epsilon\}$ (1)
 where $\mathcal{M}_{f}\in\mathbb{R}^{T\times 3},\mathcal{M}_{b}\in\mathbb{R}^{T\times 63}$ and $\mathcal{M}_{h}\in\mathbb{R}^{T\times 90}$ denotes jaw poses, body poses and hand poses, respectively.
@@
 The quantized code can be calculated as
-$Q(z^{\mathcal{M}};\mathcal{C}^{\mathcal{M}})=\mathop{\arg\min}_{k}\|z-c_{k}^{\mathcal{M}}\|_{2}.$ (3)
+$Q(z^{\mathcal{M}};\mathcal{C}^{\mathcal{M}})=\mathop{\arg\min}_{k}\|z^{\mathcal{M}}-c_{k}^{\mathcal{M}}\|_{2}.$ (3)
 The tokenizer decoder $\varphi$ will project the quantized code back into the human motion after quantization. The reconstruction objective in the tokenizer training stage can be represented as
@@
 4.3 Vocal2unit Audio Tokenizer
-Overall, we leverage the self-supervised framework polyak2021speech in speech resynthesis domain to learn vocal representations from the audio sequences.
+Overall, we leverage the self-supervised framework [polyak2021speech] in speech resynthesis domain to learn vocal representations from the audio sequences.
 Specifically, we train a Vocal2unit audio tokenizer to build a discrete tokenized representation for the human singing voice.
@@
 Semantic Encoder.
-In order to extract the semantic information of the audio into discrete tokens, we use a pre-trained Hubert encoder hsu2021hubert as the main component of our semantic encoder network $E^{h}$.
-Given a raw audio sequence $v$, the output of the Hubert encoder is continuous representation.
+In order to extract the semantic information of the audio into discrete tokens, we use a pre-trained Hubert encoder [hsu2021hubert] as the main component of our semantic encoder network $E^{h}$.
+Given a raw audio sequence $v$, the output of the Hubert encoder is a continuous representation.
 Then we apply a K-means algorithm on the representations to finally generate discrete units $z^{\mathcal{S}}=\{z^{\mathcal{S}}_{i}\}_{i=1}^{L_{s}}$ where $z^{\mathcal{S}}_{i}\in\{1,2,\cdots,K\}$.
@@
 F0 Encoder.
-We use the YAAPT algorithm kasi2002yet to extract the F0 from the input audio signal $v$, which is then fed into an encoder $E_{p}$ to generated low-frequency F0 pitch representation.
-The encoder $E^{p}$ is trained using a VQ-VAE framework with Exponential Moving Average updates similar to dhariwal2020jukebox , maintaining a learnable codebook $\mathcal{C}^{p}=\{c_{k}^{p}\}_{k=1}^{d_{c}}$, where $d_{c}$ is the number of codes in the codebook.
+We use the YAAPT algorithm [kasi2002yet] to extract the F0 from the input audio signal $v$, which is then fed into an encoder $E_{p}$ to generated low-frequency F0 pitch representation.
+The encoder $E^{p}$ is trained using a VQ-VAE framework with Exponential Moving Average updates similar to [dhariwal2020jukebox], maintaining a learnable codebook $\mathcal{C}^{p}=\{c_{k}^{p}\}_{k=1}^{d_{c}}$, where $d_{c}$ is the number of codes in the codebook.
 The raw F0 sequence is first preprocessed according to the singer, converted into a sequence of latent vectors, then mapped to its nearest neighbors in the codebook $c^{p}$, which could be represented using integer number $k\in\{1,2,\cdots,d_{c}\}$ as its index in the codebook.
-The index sequence will be further formatted as the token <pitch_k> to produce pitch token sequence $z^{\mathcal{P}}=\{z^{\mathcal{P}}_{i}\}_{i=1}^{L_{p}}$, where $L_{p}$ is the length of the pitch token sequence.
+The index sequence will be further formatted as the token ¡pitch_k¿ to produce pitch token sequence $z^{\mathcal{P}}=\{z^{\mathcal{P}}_{i}\}_{i=1}^{L_{p}}$, where $L_{p}$ is the length of the pitch token sequence.
 Finally, the decoder will reconstruct the F0 representation leveraging the pitch token sequence.
 Singer Encoder.
-As the last component of our encoders, the singer encoder $E_{si}$ from heigold2016end ; polyak2021speech is used to extract the Mel-spectrogram from the raw audio sequence and to output a singer representation $z^{\mathcal{I}}\in\mathbb{R}^{256}$.
+As the last component of our encoders, the singer encoder $E_{si}$ from [heigold2016end, polyak2021speech] is used to extract the Mel-spectrogram from the raw audio sequence and to output a singer representation $z^{\mathcal{I}}\in\mathbb{R}^{256}$.
 The singer embedding only depends on the singer and is trained globally.
 Vocoder.
-In order to decode the vocal signal from the discrete tokens, we adopt a modified version of the HiFi-GAN neural vocoder similar to the one in polyak2021speech , consisting of a generator and multiple discriminators.
+In order to decode the vocal signal from the discrete tokens, we adopt a modified version of the HiFi-GAN neural vocoder similar to the one in [polyak2021speech], consisting of a generator and multiple discriminators.
 The generator takes the encoded discrete representation $z^{\mathcal{S}}$ and $z^{\mathcal{P}}$, and the singer embedding vector $z^{\mathcal{I}}$ as inputs.
@@
 The vocal and motion tokenizers are based on VQ-VAE architectures, enabling us to represent these two modalities as discrete tokens.
-We also adopt a T5-Tokenizer raffel2020exploring to convert lyric text into tokens, so that all three modalities $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$ are united in the token space $\mathcal{T}^{\mathcal{X}}=\{t_{k}^{\mathcal{X}}\}_{k=1}^{\mathcal{N}^{\mathcal{X}}}$, where $\mathcal{N}^{\mathcal{X}}$ is the token sequence length for each modality.
+We also adopt a T5-Tokenizer [raffel2020exploring] to convert lyric text into tokens, so that all three modalities $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$ are united in the token space $\mathcal{T}^{\mathcal{X}}=\{t_{k}^{\mathcal{X}}\}_{k=1}^{\mathcal{N}^{\mathcal{X}}}$, where $\mathcal{N}^{\mathcal{X}}$ is the token sequence length for each modality.
 The large foundation model, a temporal transformer based on the decoder-only architecture, performs the next-token-prediction task based on textual features encoded by the T5-Encoder.
@@
 During the training period, we calculate the cross-entropy loss to measure the likelihood between ground truth tokens and predicted token probabilities as
-$-\sum\limits_{i=1}^{\mathcal{N^{\mathcal{X}}}}\log p_{\delta}(t^{\mathcal{X}}_{i};h^{\mathcal{L}}).$ (5)
+$\sum\limits_{i=1}^{\mathcal{N^{\mathcal{X}}}}\log p_{\delta}(t^{\mathcal{X}}_{i};h^{\mathcal{L}}).$ (5)
 After optimizing via this training objective, our model learns to predict the next token, which can be decoded into different modality features.
-This process is similar to text word generation in language models, while the “word” in our method such as <face_0123>, does not have explicit semantic information, but can be decoded into continuous modality features.
+This process is similar to text word generation in language models, while the “word” in our method such as ¡face_0123¿, does not have explicit semantic information, but can be decoded into continuous modality features.
 Multimodality Tokens Organization.
 Since we have designed a coherent vocal and motion generation pipeline with multiple token types in our method, including hubert and pitch tokens in rap vocals, and face, body, hand in whole-body motions, the organization method of all these tokens in one sentence is important for training.
-Similar to previous works lu2023humantomato ; jiang2024motiongpt , we adopt an interleaved style within each modality.
+Similar to previous works [lu2023humantomato, jiang2024motiongpt], we adopt an interleaved style within each modality.
 Specifically, with the input whole-body motion tokens $\mathcal{T}^{\mathcal{M}_{s}}=\{t^{\mathcal{M}_{s}}_{k}\}_{k=1}^{\mathcal{N}^{\mathcal{M}_{s}}}$, where $s=f,b,h$ which denotes face, body and hand respectively, and $\mathcal{N}^{\mathcal{M}_{s}}$ denotes the token length for each part.
@@
 The textual input is encoded as features to guide token inference.
-We also adopt a top-k algorithm to control the diversity of the generated content by adjusting the temperature, as generating vocals and motions based on lyrics is a creation process with multiple possible answers.
+We adopt a top-k algorithm to control the diversity of the generated content by adjusting the temperature, as generating vocals and motions based on lyrics is a creation process with multiple possible answers.
 After token prediction, a decoupling algorithm is used to process output tokens to make sure tokens from different modalities are separated and temporally aligned.
@@
 For motion synthesis, we evaluate the generation quality of the body hand gestures and the realism of the face, respectively.
-Specifically, for gesture generation, we use Frechet Inception Distance (FID) based on a feature extractor from guo2022generating to evaluate the distance of feature distributions between the generated and real motions, and Diversity (DIV) metric to assess the motions diversity.
-For face generation, we compare the vertex MSE xing2023codetalker and the vertex L1 difference LVD yi2023generating .
-Finally, we adopt Beat Constancy (BC) li2021ai to measure the synchrony of generated motion and singing vocals.
+Specifically, for gesture generation, we use Frechet Inception Distance (FID) based on a feature extractor from [guo2022generating] to evaluate the distance of feature distributions between the generated and real motions, and Diversity (DIV) metric to assess the motions diversity.
+For face generation, we compare the vertex MSE [xing2023codetalker] and the vertex L1 difference LVD [yi2023generating].
+Finally, we adopt Beat Constancy (BC) [li2021ai] to measure the synchrony of generated motion and singing vocals.
 Baselines.
-We compare the vocal generation quality with the state-of-the-art vocal generation method DiffSinger liu2022diffsinger .
-And we also adapt the text-to-speech model FastSpeech2 ren2020fastspeech for vocal generation.
+We compare the vocal generation quality with the state-of-the-art vocal generation method DiffSinger [liu2022diffsinger].
+And we also adapt the text-to-speech model FastSpeech2 [ren2020fastspeech] for vocal generation.
 For motion generation, we compare our method with both text-to-motion methods and audio-to-motion methods.
-For text-to-motion methods, since there is no existing open-sourced work for text to whole-body motion generation, we compare with transformer-based T2M-GPT zhang2023t2m and MLD chen2023executing for body generation.
+For text-to-motion methods, since there is no existing open-sourced work for text to whole-body motion generation, we compare with transformer-based T2M-GPT [zhang2023t2m] and MLD [chen2023executing] for body generation.
 For the audio-to-motion generation, we compare with Habibie et al.
-habibie2021learning and the SOTA model Talkshow yi2023generating .
+[habibie2021learning], Talkshow [yi2023generating], Motioncraft [bian2024adding] and EMAGE [liu2023emage].
 We report all the results on RapVerse with an 85%/7.5%/7.5% train/val/test split.
@@
 Text-to-Motion
-T2M-GPT zhang2023t2m 18.42 11.75 - - -
-MLD chen2023executing 24.52 12.15 - - -
+T2M-GPT [zhang2023t2m] 23.45 11.75 - - -
+MLD [chen2023executing] 26.34 12.15 - - -
 Audio-to-Motion
-Habibie et al. habibie2021learning 32.14 10.08 0.476 2.11 9.54
-Talkshow yi2023generating 21.10 13.14 0.487 1.93 9.20
-Cascaded Result 23.42 12.87 0.479 2.05 9.38
+Habibie et al. [habibie2021learning] 32.14 10.08 0.476 2.13 9.54
+Talkshow [yi2023generating] 18.23 13.14 0.482 2.05 9.20
+EMAGE [liu2023emage] 21.18 12.65 0.488 1.96 8.45
+MotionCraft [bian2024adding] 17.75 13.76 0.482 2.06 9.23
+Cascaded Result 23.42 12.87 0.479 2.09 9.38
 Text-to-Audio+Motion
-Ours 18.65 12.19 0.481 2.03 7.23
-(a) Motion Generation
+Ours 17.58 14.08 0.485 2.03 7.23
 […]
@@
 Reconstruction 4.02 ± 0.08
-FastSpeech2 ren2020fastspeech 3.41 ± 0.18
-DiffSinger liu2022diffsinger 3.72 ± 0.12
+FastSpeech2 [ren2020fastspeech] 3.41 ± 0.18
+DiffSinger [liu2022diffsinger] 3.72 ± 0.12
 Ours 3.64 ± 0.15
-(b) Vocal Generation
 Table 2: Quantitative results of generated motion and vocal.
@@
 As is demonstrated, our model rivals with both text-to-motion and audio-to-motion methods in terms of metrics measuring body motion quality and face motion accuracy.
-Furthermore, the cornerstone of our approach lies in the simultaneous generation of vocals and motion, aiming to achieve temporal alignment between these modalities.
+[…]
+GT 0 14.53 0.499 0 0
+Pretrained LLM 48.25 12.65 0.462 3.28 12.15
+Single motion token 19.15 12.75 0.477 2.18 10.12
+Ours 17.58 14.08 0.485 2.03 7.23
+Table 3: Ablation study. We compare with common designs in motion generation frameworks.
+Figure 4: User study on generated motion.
+Furthermore, the cornerstone of our approach lies in the simultaneous generation of vocals and motion, aiming to achieve temporal alignment between them.
 This objective is substantiated by our competitive results on the BC metric, which assesses the synchrony between singing vocals and corresponding motions, underscoring our success in closely synchronizing the generation of these two modalities.
@@
 Specifically, if the text-to-vocal module produces unclear vocals, it subsequently hampers the audio-to-motion model’s ability to generate accurate facial expressions that align with the vocal content.
+To better evaluate the generation quality, we conducted an additional user study (see Fig.
+4) using ten 10-second-long videos from the RapVerse dataset, sampled from the test set.
+We recruited 20 participants to evaluate four aspects: the holistic realism of human motion, expression-vocal synchronism, gesture-vocal synchronism, and holistic diversity, using real data as a reference.
+Participants were asked to sort shuffled videos generated by different methods based on these criteria.
+As shown in the results, our model achieved the highest scores in human motion realism, gesture-vocal synchronism, and holistic diversity when compared with T2MGPT (text-to-motion) and Talkshow (audio-to-motion).
 Evaluations on vocal generations.
-We have also compared our method against other state-of-the-art text-to-vocal generation baselines in Table.
+We compare our method against other state-of-the-art text-to-vocal generation baselines in Table.
 2 (b).
-While our unified model is trained to simultaneously generate vocals and motion, a task considerably more complex than generating vocals alone, its vocal generation component still manages to achieve results comparable to those systems designed solely for vocal generations.
-[…]
-GT 0 12.53 0.499 0 0
-Pretrained LLM 48.25 10.05 0.462 3.28 12.15
-Single motion token 19.15 11.29 0.477 2.18 10.12
-Ours 18.65 12.19 0.481 2.03 7.23
-Table 3: Ablation study. We compare with common designs in motion generation frameworks.
+While our unified model is trained to simultaneously generate vocals and motion, a task considerably more complex than generating vocals alone, its vocal generation component still manages to achieve results comparable to systems designed solely for vocal generations.
 5.3 Ablation Study
@@
 3.
-Initially, we explored the integration of a pre-trained large language model raffel2020exploring for multi-modality generation, akin to the approach in  jiang2024motiongpt .
+Initially, we explored the integration of a pre-trained large language model [raffel2020exploring] for multi-modality generation, akin to the approach in  [jiang2024motiongpt].
 However, the efficacy of utilizing pre-trained language models significantly lags behind our tailored design, underscoring that pre-training primarily on linguistic tokens does not facilitate effective prediction across multiple modalities like vocal and motion.
@@
 We anticipate that this work will inspire novel avenues in the joint modeling of text, audio, and motion.
-Limitations and Future Works.
-One limitation of RapVerse is that it is currently constrained to the rap music generation, excluding other music genres.
-However, we want to emphasize that our method is a general framework that could be flexibly utilized in other scenarios for joint audio and motion generation if provided with the datasets.
-Another important future direction resides in multi-performer audio and motion generation, which could be utilized in contemporary music performances such as virtual live bands.
-Appendix A Appendix
-The supplementary material is organized as follows: Sec.
-A.1 provides a website demo to show additional qualitative results; Sec.
-A.2 presents additional details of the network architectures; Sec.
-A.3 introduces evaluation metrics; Sec.
-A.4 presents additional ablation studies; Sec.
-A.5 shows additional qualitative results; Sec.
-A.6 discusses broader societal impacts of the work.
-A.1 Website Demo
-In order to provide more vivid and clear qualitative results, we make a supplemental website demo to demonstrate the generation quality of our proposed system.
-We encourage readers to view the results at https://vis-www.cs.umass.edu/RapVerse/.
-A.2 Implementation Details
-Motion Tokenizer.
-We train three separate Vector Quantized Variational Autoencoders (VQ-VAE) for face, body and hand, respectively.
-We adopt the same VQ-VAE architecture based on guo2022tm2t ; zhang2023t2m ; jiang2024motiongpt .
-For loss functions optimizing the motion tokenizers, we use the L1 smooth reconstruction loss, the embedding loss, and the commitment loss.
-The commitment loss weight is set to 0.02.
-In line with guo2022tm2t ; jiang2024motiongpt , strategies such as the exponential moving average and the codebook reset technique razavi2019generating are implemented to optimize codebook efficacy throughout the training process.
-We take 512 for the codebook size and set the dimension of each code to 512.
-We set the temporal down-sampling rate to 4.
-We train the VQ-VAEs with a batch size of 256 and a sequence window length of 72.
-We adopt Adam with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of 0.0002 as the optimizer.
-Vocal Tokenizer.
-For the semantic encoder, we adopt a BASE 12 transformer of HuBERT hsu2021hubert pre-trained on the 969-hour LibriSpeech corpus panayotov2015librispeech .
-Following lakhotia2021generative ; polyak2021speech , we derive the feature activations from its sixth layer.
-This process allows the HuBERT model to transform input audio into a 768-dimensional vector space.
-Subsequently, we employ the k-means algorithm with 500 centroids to get quantized discrete content codes.
-For the F0 encoder, a VQ-VAE framework is utilized to discretize the F0 signal into quantized F0 tokens.
-We adopt the Exponential Moving Average updates during training the VQ-VAE following dhariwal2020jukebox ; polyak2021speech .
-We set the codebook size of the VQ-VAE to 20 entries.
-Moreover, as the original work directly normalizes the extracted F0 values for each singer respectively, we don’t explicitly use the singer’s statistics but adopt a windowed convolutional layer with both the audio input (sliced into the window size) and singer’s embedding as input.
-Finally, we adopt a similar architecture as heigold2016end for the singer encoder.
-General Auto-regressive Model.
-The auto-regressive model consists of a T5 Embedder and a Foundation Model.
-We use a T5-Large Encoder as our Embedder, with 24 layers and 16 heads.
-The Embedder is freezed during foundation model training.
-The foundation model is based on the Decoder-only transformer architecture, which has 12 layers and 8 heads.
-We use Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, and a learning rate of 0.0002.
-We do not use dropout in our training.
-Our training batch size is 384 for 100 epochs.
-A.3 Evaluation Metrics
-To evaluate the motion generation quality, we utilize the following metrics:
-  1. 1.
-    Frechet Inception Distance (FID): This metric measures the distribution discrepancy between the ground truth and generated motions of body and hand gestures.
-Specifically, we train an autoencoder based on guo2022tm2t as the motion feature extractor.
-  2. 2.
-    Diversity (DIV): DIV evaluates the diversity of generated motions, where we calculate the variance from the extracted motion features.
-  3. 3.
-    Vertex MSE: Following xing2023codetalker , we compute the mean L2 error of lip vertices between generated face motions and ground truth.
-  4. 4.
-    Landmark Velocity Difference (LVD): Introduced by yi2023generating , LVD calculates the velocity difference of generated facial landmarks and ground truth.
-  5. 5.
-    Beat Constancy (BC) li2021ai : BC evaluates the synchrony of generated motions and singing vocals by calculating the similarity between the rhythm of gestures and audio beat.
-Specifically, we extract audio beats using librosa mcfee2015librosa , and we compute the kinematic beats as the local minima of the velocity of joints.
-Then the alignment score is derived from the mean proximity of motion beats to the nearest audio beat.
-For the evaluation of singing vocal generation quality, the Mean Opinion Score (MOS) is employed.
-It reflects the perceived naturalness of the synthesized vocal tones, with human evaluators rating each sample on a scale from 1 to 5, thereby offering a subjective measure of vocal synthesis fidelity.
-A.4 Additional Ablation Studies
-Ablation on Motion Tokenizer.
-We study different designs of our motion tokenizer by comparing the reconstruction results.
-Specifically, we explore VQ-VAEs with different codebook sizes, and study the effect of using a single VQ-VAE to model full-body motions instead of multiple VQ-VAEs for different body parts.
-As is demonstrated in Table.
-4, we find that using separate VQ-VAEs for face, body and hands has lower reconstruction error.
-And we select a codebook size of 512 for our final model.
-[…]
-K = 256 63.4 39.1 9.14
-K = 512 60.9 35.4 8.90
-K = 1024 62.2 36.8 9.08
-Single 65.2 38.8 9.32
-Ours 60.9 35.4 8.90
-Table 4: Evaluation of our motion tokenizer.
-We follow lu2023humantomato to evaluate the motion reconstruction errors of our motion VQVAE model $V_{m}$.
-MPJPE and PAMPJPE are measured in millimeters.
-ACCL indicates acceleration error.
-Ablation on Vocal Tokenizer.
-We also study different designs for our audio tokenizer by comparing the reconstruction results.
-Specifically, we explore different codebook sizes for the semantic encoder by changing the K-Means number.
-We also compare the effect with our singer embedding in F0 value postprocessing.
-We use the following metrics to measure the reconstruction quality of the vocal tokenizer:
-  1. 1.
-    Character Error Rate (CER): We use Whisper radford2022whisper to transcribe the ground truth and synthesized audios, and then take the corresponding ground truth lyrics as the reference to calculate the CER of the synthesized audios.
-  2. 2.
-    Gross Pitch Error (GPE): The percentage of pitch estimates that have a deviation of more than 20% around the true pitch.
-Only frames considered pitched by both the ground truth and the synthesized audio are considered.
-  3. 3.
-    Voicing Decision Error (VDE) NAKATANI2008203 : The portion of frames with voicing decision error, that is, the results using ground truth and synthesized audio to determine whether the frame is voiced are different.
-[…]
-GT 41.88 - -
-K = 100 76.39 2.64 11.98
-K = 500 (Ours) 69.21 2.29 8.84
-K = 2000 67.46 2.54 9.08
-Original F0 post-process 68.93 2.71 9.19
-Table 5: Evaluation of our unit2wav model.
-We follow polyak2021speech to evaluate the speech resynthesis errors of our unit2wav model.
-WER, GPE, and VDE, expressed as percentages, indicate the character error rate, the grand pitch error and the voicing decision error.
-The analysis of the results indicates that even the original audio exhibits a high CER, which could be attributed to the rapid speech rate associated with rapping.
-In some instances, the lyrics may not be distinctly recognizable even by human listeners.
-Upon comparing different codebook sizes, it is observed that they achieve comparable GPE values.
-This similarity in GPE is expected since the same F0 model is employed across all codebook sizes.
-The CER, which serves as a direct measure of the semantic information preserved in the code, suggests that larger codebooks tend to retain more semantic information.
-However, the difference in CER between codebook sizes of $K=500$ and $K=2000$ is minimal.
-Given that $K=500$ demonstrates better GPE and VDE, we select $K=500$.
-Additionally, we ablate a design without the singer embedding in the F0 preprocessing, instead normalizing the F0 values for each singer.
-It shows that this approach resulted in significantly inferior performance, particularly in pitch prediction, compared to the modified version that includes the singer embedding.
-A.5 Additional Qualitative Results
-We show additional qualitative results in Fig.
-4.
-Our model adeptly generates comprehensive whole-body motions that embody the essence of the input lyrics.
-These include authentic gesture movements that resonate with the song’s rhythm and synchronized lip motions that articulate the lyrics.
-Figure 4: Additional qualitative results. Our method can generate diverse whole-body motions from input lyrics.
-A.6 Broader Impacts
-This research contributes to advancements in generating synchronized vocals and human motion from textual lyrics, aiming to enhance virtual agents’ ability to provide immersive and interactive experiences in digital media.
-The potential positive impact of this work lies in its ability to create more lifelike and engaging virtual performances, such as in virtual concerts and gaming, where characters can perform and react in ways that are deeply resonant with human expressions.
-This could significantly enhance user engagement in virtual reality settings, and provide innovative solutions in entertainment industries.
-However, this capability carries inherent risks of misuse.
-The technology’s ability to generate realistic human-like actions and singing vocals from mere text raises concerns about its potential to create misleading or deceptive content.
-For example, this could be exploited to produce fake videos or deepfakes where individuals appear to sing and perform that never actually occurred, which could be used to spread misinformation or harm reputations.
-Recognizing these risks, it is crucial to advocate for ethical guidelines and robust frameworks to ensure the responsible use of such technologies.